# Molecule Language Model with Augmented Pairs and Expertise Transfer

### v1

@@ 4.2 Structural Similarity Preserving Loss
While we increase the number of pairs by sharing the descriptions, training the model with augmented pairs is challenging since the shared description $t_{i}$ is not specifically written for the substituted molecule $\mathcal{G}_{i^{\prime}}$. Consequently, employing traditional InfoNCE loss in Eq. 2 for model training could make the model prone to false positives.
@@
$y_{ij^{\prime}}^{t\rightarrow m}=\frac{\exp({s_{ij^{\prime}}}/\tau_{1})}{\sum_{k^{\prime}=1}^{N_{\text{batch}}}\exp({s_{ik^{\prime}}}/\tau_{1})},$ (3)
@@
$\hat{y}_{ij^{\prime}}^{t\rightarrow m}=\frac{\exp(\texttt{sim}(z_{t_{i}},z_{\mathcal{G}_{j^{\prime}}})/\tau_{2})}{\sum_{k^{\prime}=1}^{N_{\text{batch}}}\exp(\texttt{sim}(z_{t_{i}},z_{{\mathcal{G}}_{k^{\prime}}})/\tau_{2})},$ (4)
@@
where $\tau_{2}$ is a temperature hyperparameter. It’s worth noting that in order to approximate the representation similarity $\texttt{sim}(z_{t_{i}},z_{\mathcal{G}_{j^{\prime}}})$ between a molecule and text to structural similarity $s_{ij^{\prime}}$, we generate pseudo labels by applying softmax normalization to structural similarity. Then AMOLE is trained with the following cross-entropy loss:
$\mathcal{L}^{t\rightarrow m}_{S^{2}P}=-\frac{1}{N_{Batch}}\sum_{i=1}^{N_{batch}}\sum_{j^{\prime}=1}^{N_{batch}}{y_{ij^{\prime}}^{t\rightarrow m}\log\hat{y}_{ij^{\prime}}^{t\rightarrow m}}.$ (5)
@@
By doing so, the structural similarity between molecules $\mathcal{G}_{i}$ and $\mathcal{G}_{i^{\prime}}$ instructs the model on how to align the text $t_{i}$ and molecule $\mathcal{G}_{i^{\prime}}$ closely in the representation space, enabling the selective sharing of descriptions based on structural similarities. For the symmetricity, we compute $\mathcal{L}^{m\rightarrow t}_{S^{2}P}$ and derive our final training loss: $\mathcal{L}_{S^{2}P}=\mathcal{L}^{t\rightarrow m}_{S^{2}P}+\mathcal{L}^{m\rightarrow t}_{S^{2}P}$.
4.3 Expertise Transfer Module
@@
Experts in the field often have a specific area of focus, which results in some molecules lacking comprehensive expertise coverage from different specialists. To address this issue of missing expertise, we suggest transferring the expertise gained from molecules with extensive descriptions to those with less description. Specifically, we use a molecule $\mathcal{G}_{i}$ with a set of descriptions $\mathcal{T}_{i}$ obtained from various experts. We train the language model $f_{\text{text}}$ to reconstruct the description $t_{i^{\prime}}\in\mathcal{T}_{i}$ from a given description $t_{i}$. This approach allows the language model $f_{\text{text}}$ to become skilled at inferring expertise from one institution based on another. It behaves as if it has access to abundant expertise even when dealing with molecules with missing expertise.
@@
However, reconstructing the description $t_{i^{\prime}}$ in text space poses a challenge, as our model is designed to learn a qualified representation space without a decoder structure. To this end, we propose a novel expertise reconstruction ($ER$) loss, which guides the model to reconstruct the description within the representation space rather than directly in the text space. More formally, when we have a textual description $t_{i}$ and another associated description $t_{i^{\prime}}$ for a molecule $\mathcal{G}_{i}$, we formulate the reconstruction target $\tilde{t}_{i}$ by concatenating the two descriptions with [SEP] token, i.e., $\tilde{t}_{i}=$ $t_{i}$ [SEP] $t_{i^{\prime}}$. Then, the model is trained to minimize the L2 distance as follows:
$\mathcal{L}_{ER}=-\frac{1}{N_{Batch}}\sum_{i=1}^{N_{batch}}{\|f_{\text{text}}(t_{i})-\text{SG}(f_{\text{text}}(\tilde{t}_{i}))\|^{2}_{2}},$ (6)
@@
where SG denotes the stop-gradient operation, which halts the propagation of gradients when inputs are structured as $\tilde{t_{i}}$, thereby preventing the model from acquiring degenerate solutions by disregarding the text following the [SEP] token. With the novel $ER$ training loss, the model simulates having extensive expertise available, even when specific expertise is lacking.
4.4 Model Training
@@
$\mathcal{L}=\mathcal{L}_{S^{2}P}+\alpha\cdot\mathcal{L}_{ER},$ (7)
@@
5 Experiments
5.1 Experimental Setup
Pretraining Dataset. We pre-train AMOLE with PubChem database Kim et al. (2021), which is one of the most extensive public molecular databases available. Our pre-training dataset is compiled using the preprocessing script provided in the repository of the previous work Liu et al. (2023a). However, due to regular updates to the PubChem database, our dataset varies slightly from the previous work, comprising a total of 299K unique molecules and 336K molecule-text pairs. We provide further details on PubChem database and diverse datasets utilized across four distinct downstream tasks in Appendix C.
@@ 5.1 Experimental Setup
Table 1: Model accuracy (%) in zero-shot cross-modal retrieval task. The value within the brackets indicates the variance observed across five trials.
@@
Table 2: Ablation studies results.
@@
Table 3: Model accuracy (%) in zero-shot question and answering task results.
Implementation Details. One key benefit of treating molecules and text as distinct modalities is the opportunity to leverage powerful pre-trained models specifically designed for each modality. Following a previous work Liu et al. (2023a), we employ a GraphMVP Liu et al. (2021) pre-trained checkpoint for graph isomorphism network (GIN) model Xu et al. (2018) for our molecule encoder $f_{\text{mol}}$. Moreover, for our text encoder $f_{\text{text}}$, we utilize a pre-trained SciBERT Beltagy et al. (2019), which has been trained on a vast corpus of textual data from the biochemistry domains. We provide further implementation details in Appendix A.
Baseline Methods. We evaluate AMOLE against three single encoder models: MolT5 Edwards et al. (2022), BioT5 Pei et al. (2023), and KV-PLM Zeng et al. (2022), all of which represent molecules as 1D SMILES strings. While T5-based models are not designed for learning representations of molecules and text, we assess their capabilities by examining the hidden representations produced by the encoder model following a previous work Seidl et al. (2023). Additionally, we compare AMOLE with two separate encoder models: MoMu Su et al. (2022) and MoleculeSTM Liu et al. (2023a). While MoMu depicts molecules as 2D graph structures, MoleculeSTM introduces two models that utilize both SMILES and 2D graph representations for molecules. For the single encoder models, we utilize the checkpoints provided by the authors of the original papers. However, for models with separate encoders, we independently pre-train models with the same pre-training data and model architecture for fair comparison. We provide further details on baseline methods in Appendix B.
5.2 Zero-Shot Cross-Modal Retrieval
Task Description. This task requires choosing an appropriate description from several alternatives for a specific molecule (Given Molecule) or retrieving the molecule that aligns with a particular description (Given Text). Following a previous work Liu et al. (2023a), our experiments are carried out with varying numbers of choices, specifically 4, 10, and 20 options. We report the models’ performance on 20 options in Table 3 of the main manuscript, while the complete set of experimental results is available in the Appendix E.1. We also provide details on the datasets and evaluation scheme in Appendix C.2 and D.1, respectively.
Empirical Results. In Table 3, we have the following observations: 1) As T5-based models (i.e., MolT5 and BioT5) are not intended to build joint representation space between molecule and text, their performance was the worst among the models. This suggests that T5-based models struggle to learn suitable representations and necessitate a costly fine-tuning process after the pre-training step. 2) Additionally, despite the SMILES-based MoleculeSTM having a significantly higher number of parameters compared to the graph-based model,  [^1] For MoleculeSTM SMILES, we use ChemBERTa, which contains 83,450,880 parameters, and for MoleculeSTM Graph, we use GIN, which contains 1,885,206 parameters. SMILES and graph structure representations of molecules yield comparable results. This underscores the efficacy of graph-based representations in capturing molecular properties. 3) Within the graph-based models, it is evident that MoMu generally exhibits the lowest performance across all but one dataset. This is largely due to the random augmentation of molecules, such as by dropping nodes, without taking the chemical validity of the molecules into account Lee et al. (2022b), which makes the model perform even worse than MoleculeSTM. On the other hand, in the upcoming Ablation studies, we will demonstrate that augmenting molecule-text pairs with chemically valid molecules can consistently improve MoleculeSTM performance. 4) Overall, we observe that AMOLE performs the best on five out of six datasets, demonstrating its ability to comprehend and bridge two heterogeneous modalities.
@@ 5.2 Zero-Shot Cross-Modal Retrieval
Table 4: ROC-AUC performance in molecular property prediction task. The value within the brackets indicates the variance observed across five trials. Bold text denotes the top performance, while an underline highlights the second best. ”Avg. Rank" represents the average ranking across all datasets.
Ablation Studies. Now, we empirically evaluate the impact of each component in AMOLE by sequentially removing them one at a time. We assess the effectiveness of each component by calculating the average performance across hard cross-retrieval tasks, including retrieving among 20 texts given a molecule and retrieving among 20 molecules given a text. In Table 3, we have the following observations: 1) By comparing MoleculeSTM and Ablation 1, we observe that augmenting molecule-text consistently brings performance gain, indicating that “scale-is-everything” perspective might also be applicable and beneficial within the MoLM community. 2) Additionally, a comparison between Ablation 1 and Ablation 2 reveals that incorporating a $S^{2}P$ loss can further enhance the performance. This indicates that aligning molecules and text based on the structural similarity among molecules can effectively mitigate issues related to false positives. 3) Lastly, it is noted that the $ER$ loss uniformly enhances performance across all datasets, with a particularly notable improvement in the ATC dataset. This improvement can be attributed to the nature of the ATC dataset, which comprises labels for a molecule classification system and is inherently abstract. Consequently, reconstructing missing expertise from such abstract descriptions proves advantageous for model performance. We provide further model analysis and statistical significance tests in Appendix E.1.
In summary, our model effectively augments molecule-text pair by incorporating chemically viable molecules and employing a $S^{2}P$ loss. Additionally, transferring the knowledge through $ER$ loss consistently enhances AMOLE performance, particularly with brief and abstract descriptions.
5.3 Zero-Shot Question and Answering
@@
Task Description. While earlier studies have assessed the retrieval capabilities of MoLM by contrasting the correct description with randomly chosen descriptions from other molecules Liu et al. (2023a); Su et al. (2022), we propose a novel task, termed the “Zero-shot Question and Answering” task. Specifically, given a textual description of a molecule, we instruct GPT-4 Achiam et al. (2023) to generate a multiple-choice question comprising five options, all derived from the given textual description. Then, with a generated question and its five corresponding options, we merge the question with each option to form a single input, i.e., $\texttt{input}_{i}=\text{Concat}(\texttt{question},~{}\texttt{option}_{i})$, where $i=1,\ldots,5$. Given a molecule and these combined inputs, we then select the one that includes the correct answer from the options. In Appendix C.3, we offer details for generating and validating questions and answer datasets, along with an evaluation scheme for the task in Appendix D.2.
@@
Empirical Results. In Table 3, we have the following observations: 1) Comparing to Table 3, we observe that despite having far fewer options for retrieval, most models learning representation space (i.e., KV-PLM, MoMu, MoleculeSTM, and AMOLE) perform much worse in the task. This is because the model must discern based solely on the subtle differences between the options provided, requiring the model to have a more fine-grained understanding of molecules than cross-modal retrieval. 2) On the other hand, AMOLE consistently outperforms baseline models in this task, showcasing its ability to extract more intricate information from the slight variations in options through the inference of related expertise. In conclusion, we posit that AMOLE offers benefits in tasks that demand a more intricate understanding of molecules, achieved by integrating additional expertise. In Appendix E.2, we further showcase the effectiveness of the expertise transfer module by integrating modified inputs as AMOLE into the baseline methods.
Dataset Prompt

### v3

@@ 4.2. Structural Similarity Preserving Loss
While we increase the number of pairs by sharing the descriptions, training the model with augmented pairs is challenging since the shared description $t_{i}$ is not specifically written for the substituted molecule $\mathcal{G}_{i^{\prime}}$. Consequently, employing traditional InfoNCE loss in Eq. 2, which strictly encourages the positive pairs to be close in the representation space, might not be suitable for training as the augmented pairs may not truly be positive.
@@
(3) $y_{ij^{\prime}}^{t\rightarrow m}=\frac{\exp({s_{ij^{\prime}}}/\tau_{1})}{\sum_{k^{\prime}=1}^{N_{\text{batch}}}\exp({s_{ik^{\prime}}}/\tau_{1})},$
@@
(4) $\hat{y}_{ij^{\prime}}^{t\rightarrow m}=\frac{\exp(\texttt{sim}(z_{t_{i}},z_{\mathcal{G}_{j^{\prime}}})/\tau_{2})}{\sum_{k^{\prime}=1}^{N_{\text{batch}}}\exp(\texttt{sim}(z_{t_{i}},z_{{\mathcal{G}}_{k^{\prime}}})/\tau_{2})},$
@@
where $\tau_{2}$ is a temperature hyperparameter. We generate pseudo labels by applying softmax normalization to structural similarity since the loss aims to approximate the representation similarity $\texttt{sim}(z_{t_{i}},z_{\mathcal{G}_{j^{\prime}}})$ between a molecule and text to structural similarity $s_{ij^{\prime}}$. Then AMOLE is trained with the following cross-entropy loss:
(5) $\mathcal{L}^{t\rightarrow m}_{S^{2}P}=-\frac{1}{N_{Batch}}\sum_{i=1}^{N_{batch}}\sum_{j^{\prime}=1}^{N_{batch}}{y_{ij^{\prime}}^{t\rightarrow m}\log\hat{y}_{ij^{\prime}}^{t\rightarrow m}}.$
@@
By doing so, the structural similarity between molecules $\mathcal{G}_{i}$ and $\mathcal{G}_{i^{\prime}}$ instructs the model on how to align the text $t_{i}$ and molecule $\mathcal{G}_{i^{\prime}}$ closely in the representation space, enabling the selective sharing of descriptions based on structural similarities. For the symmetricity, we compute $\mathcal{L}^{m\rightarrow t}_{S^{2}P}$ and derive final structural similarity preserving loss: $\mathcal{L}_{S^{2}P}=\mathcal{L}^{t\rightarrow m}_{S^{2}P}+\mathcal{L}^{m\rightarrow t}_{S^{2}P}$.
4.3. Expertise Transfer Module
@@
On the other hand, experts in molecular science often have a specific area of focus, resulting in some molecules lacking comprehensive expertise from different specialists. To address this issue of missing expertise, we suggest transferring the expertise gained from molecules with extensive descriptions to those with less description. Specifically, we use a molecule $\mathcal{G}_{i}$ with a set of descriptions $\mathcal{T}_{i}$ obtained from various experts. We train the language model $f_{\text{text}}$ to reconstruct the description $t_{i^{\prime}}\in\mathcal{T}_{i}$ from a given description $t_{i}$. This approach allows the language model $f_{\text{text}}$ to become skilled at inferring expertise from one institution based on another. As a result, it behaves as if it has access to abundant expertise even when dealing with molecules with missing expertise.
@@
However, reconstructing the description $t_{i^{\prime}}$ in text space poses a challenge, as our model is designed to learn a qualified representation space without a decoder structure. To this end, we propose a novel expertise reconstruction ($ER$) loss, which guides the model to reconstruct the description within the representation space rather than directly in the text space. More formally, when we have a textual description $t_{i}$ and another associated description $t_{i^{\prime}}$ for a molecule $\mathcal{G}_{i}$, we formulate the reconstruction target $\tilde{t}_{i}$ by concatenating the two descriptions with [SEP] token, i.e., $\tilde{t}_{i}=$ $t_{i}$ [SEP] $t_{i^{\prime}}$. Then, the model is trained to minimize the L2 distance between the textual description $t_{i}$ and reconstructed target $\tilde{t}_{i}$ as follows:
(6) $\mathcal{L}_{ER}=-\frac{1}{N_{Batch}}\sum_{i=1}^{N_{batch}}{\|f_{\text{text}}(t_{i})-\text{SG}(f_{\text{text}}(\tilde{t}_{i}))\|^{2}_{2}},$
@@
where SG denotes the stop-gradient operation, which halts the propagation of gradients when inputs are structured as $\tilde{t_{i}}$, thereby preventing the model from acquiring degenerate solutions by disregarding the text following the [SEP] token. It is also worth noting that since $f_{\text{text}}$ is a language model, it can accommodate texts of varying lengths. Moreover, as illustrated in Figure 1 (b), molecules described by more than two texts are rare, and this scarcity could reduce the effectiveness of the $ER$ Loss. To mitigate this problem, we sample the $N_{batch}$ number of molecules that have more than two descriptions and apply the $ER$ loss to those molecules. In conclusion, with the training loss, the model behaves as if there exists extensive expertise available, enabling reliable predictions even when a detailed description of the molecule is lacking.
4.4. Model Training
@@
(7) $\mathcal{L}=\mathcal{L}_{S^{2}P}+\alpha\cdot\mathcal{L}_{ER},$
@@
5. Experiments
5.1. Experimental Setup
Pretraining Dataset. We pre-train AMOLE with PubChem database (Kim et al., 2021), which is one of the most extensive public molecular databases available. PubChem database consists of multiple data sources including DrugBank (Wishart et al., 2018), CTD (Davis et al., 2021), PharmGKB (Thorn et al., 2013), and more.  [^1] Please refer to the following URL for more details on the data sources included in PubChem database: https://pubchem.ncbi.nlm.nih.gov/sources/. Our pre-training dataset is compiled using the preprocessing script provided in the repository of the previous work (Liu et al., 2023b). However, due to regular updates to the PubChem database, our dataset varies slightly from the previous work, comprising a total of 299K unique molecules and 336K molecule-text pairs. During preprocessing, we consolidate each expertise into a unified description, meaning that each description of an individual molecule originates from distinct expertise. As shown in Figure 1 (b), it is evident that the majority of molecules have a singular description, indicating a lack of comprehensive expertise across various experts for numerous molecules. On average, molecules are associated with 1.115 descriptions, with a maximum of 17 descriptions and a minimum of one. Moreover, each description in training data consists of 17.62 words on average, with a maximum of 874 words and a minimum of one.
Downstream Tasks. After training the model using the PubChem database, we assess its performance on four distinct downstream tasks: 1) Zero-Shot Cross-Modal Retrieval, 2) Zero-Shot Question and Answering, 3) Molecular Property Prediction, and 4) Zero-Shot Virtual Screening. While tasks such as 1) Zero-Shot Cross-Modal Retrieval and 3) Molecular Property Prediction are commonly explored in previous studies (Liu et al., 2023b), we propose novel tasks in this paper: 2) Zero-Shot Question and Answering, and 4) Zero-Shot Virtual Screening tasks that are essential for evaluating the model’s performance on fine-grained knowledge extraction and potential for application in real-world drug discovery, respectively.
Implementation Details. Following previous work (Liu et al., 2023b), we use graph isomorphism network (GIN) (Xu et al., 2018) architecture as a molecular encoder $f_{\text{mol}}$, which has been widely used as the backbone model in recent graph self-supervised learning works (Hu et al., 2019; Liu et al., 2021). Moreover, we use BERT architecture (Devlin et al., 2018) for the text encoder $f_{\text{text}}$. One key benefit of treating molecules and text as distinct modalities is the opportunity to leverage powerful pre-trained models specifically designed for each modality. Following a previous work (Liu et al., 2023b), we employ a GraphMVP (Liu et al., 2021) pre-trained checkpoint for GIN model (Xu et al., 2018) for our molecule encoder $f_{\text{mol}}$. Moreover, for our text encoder $f_{\text{text}}$, we utilize a pre-trained SciBERT (Beltagy et al., 2019), which has been trained on a vast corpus of textual data from the biochemistry domains.
Baseline Methods. We evaluate AMOLE against three single encoder models: MolT5 (Edwards et al., 2022), BioT5 (Pei et al., 2023), and KV-PLM (Zeng et al., 2022), all of which represent molecules as 1D SMILES strings. While T5-based models are not designed for learning representations of molecules and text, we assess their capabilities by examining the hidden representations produced by the encoder model following a previous work (Seidl et al., 2023). Additionally, we compare AMOLE with two separate encoder models: MoMu (Su et al., 2022) and MoleculeSTM (Liu et al., 2023b). While MoMu depicts molecules as 2D graph structures, MoleculeSTM introduces two models that utilize both SMILES and 2D graph representations for molecules. For the single encoder models, we utilize the checkpoints provided by the authors of the original papers. However, for models with separate encoders, we independently pre-train models with the same pre-training data and model architecture for a fair comparison with the proposed AMOLE.
5.2. Zero-Shot Cross-Modal Retrieval
Table 1. Model accuracy (%) in zero-shot cross-modal retrieval task. The value within the brackets indicates the variance observed across five trials. Bold denotes the best performance.
@@ 5.2. Zero-Shot Cross-Modal Retrieval
Task Description. This task requires choosing an appropriate description from several alternatives for a specific molecule (Given Molecule) or retrieving the molecule that aligns with a particular description (Given Text). Following a previous work (Liu et al., 2023b), our experiments are carried out with 20 options, where one of the options is the matching counterpart, while the others are randomly selected from the dataset. Then, the model performance is determined by its capacity to pinpoint the correct counterpart from the options provided, such as correctly matching the description to the given molecule or vice versa. For the evaluation, we conducted five separate experiments, each featuring a distinct random selection of options, and we present both the mean and standard deviation of these experiments. Moreover, we ensure that molecules appearing in the training dataset with identical canonical SMILES are excluded to avoid data leakage. We provide further details on the datasets used for evaluation in Appendix C.1.
Empirical Results. In Table 1, we have the following observations: 1) As T5-based models (i.e., MolT5 and BioT5) are not intended to build joint representation space between molecule and text, their performance was the worst among the models. This suggests that T5-based models struggle to learn suitable representations and necessitate a costly fine-tuning process after the pre-training step. 2) Additionally, despite the SMILES-based MoleculeSTM having a significantly higher number of parameters compared to the graph-based model,  [^2] For MoleculeSTM SMILES, we use ChemBERTa, which contains 83,450,880 parameters, and for MoleculeSTM Graph, we use GIN, which contains 1,885,206 parameters. SMILES and graph structure representations of molecules yield comparable results. This underscores the efficacy of graph-based representations in capturing molecular properties. 3) Within the graph-based models, it is evident that MoMu generally exhibits the lowest performance across all but one dataset. This is largely due to the random augmentation of molecules, such as by dropping nodes, without taking the chemical validity of the molecules into account (Lee et al., 2022b), which makes the model perform even worse than MoleculeSTM. On the other hand, in the upcoming Ablation studies, we will demonstrate that augmenting molecule-text pairs with chemically valid molecules can consistently improve MoleculeSTM performance. 4) Overall, we find that AMOLE achieves the best results in five out of six datasets, illustrating its capability to understand and integrate two distinct modalities. Notably, the performance gap widens with increasingly challenging tasks (i.e., Pharma. and ATC, where overall model performance is lower), highlighting AMOLE’s intricate grasp of the modalities.
Table 2. Ablation studies results.
@@
Figure 3. Sensitivity analysis on $k$.
Ablation Studies. Now, we empirically evaluate the impact of each component in AMOLE by sequentially removing them one at a time. We assess the effectiveness of each component by calculating the average performance across cross-retrieval tasks, including retrieving among 20 texts given a molecule and retrieving among 20 molecules given a text. In Table 2, we have the following observations: 1) By comparing MoleculeSTM and Ablation 1, we observe that augmenting molecule-text consistently brings performance gain, indicating that “scale-is-everything” perspective might also be applicable and beneficial within the MoLM community. 2) Additionally, a comparison between Ablation 1 and Ablation 2 reveals that incorporating a $S^{2}P$ loss can further enhance the performance. This indicates that aligning molecules and text based on the structural similarity among molecules can effectively mitigate issues related to false positives. 3) Lastly, it is noted that the $ER$ loss uniformly enhances performance across all datasets, with a particularly notable improvement in the ATC dataset. This improvement can be attributed to the nature of the ATC dataset, which comprises labels for a molecule classification system and is inherently abstract. Consequently, reconstructing missing expertise from such abstract descriptions proves advantageous for model performance.
In summary, our model effectively augments molecule-text pair by incorporating chemically viable molecules and employing a $S^{2}P$ loss. Additionally, transferring the knowledge through $ER$ loss consistently enhances AMOLE performance, particularly with brief and abstract descriptions. Furthermore, we note the model’s robust performance across different levels of $k$, reducing the need for extensive trial-and-error in hyperparameter tuning.
5.3. Zero-Shot Question and Answering
Table 3. Model accuracy (%) in zero-shot question and answering task.
@@ 5.3. Zero-Shot Question and Answering
Table 4. Model accuracy (%) after adopting AMOLE training strategy to MolCA.
Given Molecule @ 20 Given Text @ 20 Q&A
Descr. Pharma. ATC Descr. Pharma. ATC Descr. Pharma.
MolCA 97.93 83.71 59.01 98.06 83.50 55.86 41.93 33.09
+ AMOLE 98.46 85.45 59.34 98.51 84.93 57.53 43.25 34.43
Table 5. ROC-AUC performance in molecular property prediction task. The value within the brackets indicates the variance observed across five trials. Bold text denotes the top performance, while an underline highlights the second best. ”Avg. Rank” represents the average ranking across all datasets.
@@
Task Description. While earlier studies have assessed the retrieval capabilities of MoLM by contrasting the correct description with randomly chosen descriptions from other molecules (Liu et al., 2023b; Su et al., 2022), we propose a novel task, termed the “Zero-shot Question and Answering” task. Specifically, given a textual description of a molecule, we instruct GPT-4 (Achiam et al., 2023) to generate a multiple-choice question comprising five options, all derived from the given textual description. Then, with a generated question and its five corresponding options, we merge the question with each option to form a single input, i.e., $\texttt{input}_{i}=\text{Concat}(\texttt{question},~{}\texttt{option}_{i})$, where $i=1,\ldots,5$. Given a molecule and these combined inputs, we then select the one that includes the correct answer from the options. We generate questions based on the textual descriptions used for the cross-modal retrieval task in Section 5.2. Since the ATC dataset consists of brief labels for molecules, we restrict our question generation to descriptions and pharmacodynamics datasets. Note that the only difference between the options comes from $\texttt{option}_{i}$, which makes the task much harder compared to the cross-modal retrieval task in Section 5.2. In Appendix C.2, we offer details for generating processes and statistics for questions and answer datasets.
@@
Empirical Results. In Table 3, we have the following observations: 1) Comparing to Table 1, we observe that despite having far fewer options for retrieval, most models learning representation space (i.e., KV-PLM, MoMu, MoleculeSTM, and AMOLE) perform much worse in the task. This is because the model must discern based solely on the subtle differences between the options provided, requiring the model to have a more fine-grained understanding of molecules than cross-modal retrieval. 2) On the other hand, AMOLE consistently outperforms baseline models in this task, showcasing its ability to extract more intricate information from the slight variations in options through the inference of related expertise. In conclusion, we posit that AMOLE offers benefits in tasks that demand a more intricate understanding of molecules, achieved by integrating additional expertise.
Adaptability to Various Model Architectures. In this section, we examine whether the training strategy of AMOLE can be adapted to various model architectures beyond CLIP style architecture (Radford et al., 2021), i.e., MolCA (Liu et al., 2023a). MolCA proposes to train Q-Former with three loss terms, i.e., molecule-text contrasting, molecule-text matching, and molecule captioning. To implement our training strategy, we first increase the number of molecule-text pairs with the proposed sharing strategy. Then, we replace the molecule-text contrasting loss with a structural similarity preserving ($S^{2}P$) loss, and also incorporate expertise reconstruction ($ER$) loss along with the proposed three losses. In Table 4, we observe consistent performance enhancements of the Q-Former architecture when trained with the proposed strategy, highlighting AMOLE’s broad applicability beyond CLIP-style architecture.
5.4. Molecular Property Prediction
Task Description. In this task, we assess the potential benefits of incorporating external knowledge, i.e., textual descriptions, into the molecule encoder $f_{\text{mol}}$ as done in AMOLE, in enhancing the prediction of molecular properties. We mainly compare to recent graph self-supervised learning (GraphSSL) approaches, i.e., AttrMask (Hu et al., 2019), ContextPred (Hu et al., 2019), GPT-GNN (Hu et al., 2020), InfoGraph (Sun et al., 2019), GraphMVP (Liu et al., 2021), Mole-BERT (Xia et al., 2022), and MoLMs which represent molecules with graph structure (MoLM w/ Graph). Following previous works (Liu et al., 2021, 2023b), we pre-train the molecule encoder $f_{\text{mol}}$ using each of the proposed strategies with the same pre-taining data and fine-tuning on MoleculeNet benchmark (Wu et al., 2018). During the fine-tuning stage, we initially divide the fine-tuning dataset according to scaffold information using an 8:1:1 ratio for the training, validation, and test sets, respectively. Subsequently, we fine-tune the molecular encoder $f_{\text{mol}}$ using the training data across 100 epochs. Following previous work (Liu et al., 2023b), the model’s performance is then evaluated on the test set where the hyperparameters achieve optimal performance on the validation set. We run five individual experiments and report the average and standard deviation of the results in Table 5. We provide details on the MoleculeNet dataset in Appendix C.3.
Empirical Results. In Table 5, we have following observations: 1) We observe that incorporating external textual descriptions into the pre-training phase generally enhances the prediction of molecular properties, as evidenced by improved overall performance (i.e., averaged rank across all eight tasks). This improvement is credited to the implicit influence of external domain knowledge, i.e., the textual descriptions of molecules, which is proved to be beneficial for property prediction. (Liu et al., 2023b). 2) Among the MoLM, AMOLE outperforms baseline methods on six out of eight tasks, demonstrating that the knowledge can be more efficiently transferred to molecular property prediction through our strategy. In summary, our methodology is advantageous not just for cross-modal tasks, but also for tasks within a single modality, demonstrating the versatility of AMOLE across a wide range of downstream applications.
5.5. Zero-Shot Virtual Screening
Table 6. Utilized abstractive and detailed textual descriptions for each dataset.
Dataset Textual Description